# Mutual Information Density of Massive MIMO Systems over Rayleigh-Product Channels

### 05 Jan 2024 v4

@@ C-B1 The evaluation of $U_{1}$
$\mathbb{E}\varepsilon_{1}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=\mathbb{E}\varepsilon_{1}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\mathrm{Cov}(\varepsilon_{1},\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}),$ (102)
and $\mathrm{Cov}(\varepsilon_{1},\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}})\leq\mathrm{Var}^{\frac{1}{2}}(\varepsilon_{1})\mathrm{Var}^{\frac{1}{2}}(\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}})={\mathcal{O}}(\frac{u\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}\mathcal{P}_{2}(\frac{1}{z})}{zM^{\frac{3}{2}}})$. By the integration by parts formula, we have
$\mathbb{E}\varepsilon_{1}=-\frac{\jmath un}{\sqrt{M^{3}n}}\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}=-\frac{\jmath un}{\sqrt{M^{3}n}}(\frac{\mathbb{E}\operatorname{Tr}{\mathbb{A}}\mathbb{E}\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}}{M}$ (103)
$-\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}-\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}}{M})$
@@
Noticing that $\operatorname{Tr}{\mathbb{A}}=\operatorname{Tr}({\mathbb{I}}_{M}-\frac{{\mathbb{X}}{\mathbb{X}}^{H}}{n})=0$, we have $E_{1}=0$. By (60) and $\operatorname{Tr}{\mathbb{A}}^{2}={\mathcal{O}}(M^{2})$, we have
$|E_{2}|=\frac{un}{\sqrt{M^{3}n}}|\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}+\mathrm{Cov}(\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z),\frac{\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M})|$ (104)
$\leq\frac{Kun}{\sqrt{M^{3}n}}\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{z^{2}M^{\frac{3}{2}}}={\mathcal{O}}(\frac{u\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{z^{2}n^{\frac{5}{2}}}).$
@@
$|E_{3}|\leq\frac{un}{\sqrt{M^{3}n}}\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{z^{2}M}={\mathcal{O}}(\frac{u\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{z^{2}n^{\frac{5}{2}}}).$ (105)
A similar analysis can be performed to obtain $\mathbb{E}\varepsilon_{2}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}={\mathcal{O}}(\frac{u\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{z^{2}n^{\frac{5}{2}}})$ and $\mathbb{E}\varepsilon_{3}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=\mathbb{E}\varepsilon_{3}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(\frac{\mathcal{P}_{3}(\frac{1}{z})(1+\sqrt{\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}})}{M^{2}z^{2}})$ in the last line of (C-B1). Now we turn to evaluate $\mathbb{E}\varepsilon_{3}$, which is performed as
$\mathbb{E}\varepsilon_{3}=-\mathbb{E}\sum_{i,j}\frac{u^{2}}{2M}\frac{\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}]_{j,i}=-\frac{u^{2}n}{Mn}[\frac{1}{M}\mathbb{E}\operatorname{Tr}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)$ (106)
$-\frac{1}{M}\mathbb{E}\operatorname{Tr}{\mathbb{H}}{\mathbb{H}}^{H}({\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})^{2}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)+\frac{\sigma^{2}}{2M}\operatorname{Tr}{\mathbb{X}}{\mathbb{X}}^{H}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}$
$-\frac{\sigma^{2}\operatorname{Tr}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{X}}{\mathbb{X}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)}{M}-\frac{\sigma^{2}\operatorname{Tr}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{X}}{\mathbb{X}}^{H}{\mathbb{Q}}^{2}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)}{M}$
@@
Similarly, we can show that $E_{3,i},i=2,...,3$ are all ${\mathcal{O}}(\frac{1}{Mz})$ terms. We can evaluate $\mathbb{E}E_{3,4}$ as
$\frac{2nu^{2}}{Mn}\mathbb{E}E_{3,4}$ $=-\frac{2nu^{2}}{Mn}\mathbb{E}\frac{\sigma^{2}\operatorname{Tr}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{X}}{\mathbb{X}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)}{Mn}=\frac{2nu^{2}}{Mn}[\mathbb{E}\frac{\sigma^{2}\operatorname{Tr}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)}{M}$ (108)
$-\mathbb{E}\frac{\sigma^{2}\operatorname{Tr}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)}{M}]={\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})u^{2}(1+\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{5}{2}}})}{zM}).$
Therefore, $\mathbb{E}\varepsilon_{3}={\mathcal{O}}(\frac{u^{2}\mathcal{P}(\frac{1}{z})(1+\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{5}{2}}})}{Mz})={\mathcal{O}}(\frac{u^{2}\mathcal{P}(\frac{1}{z})}{zM})$. According to the derivation in (106) to (108), we can also show that
$\mathbb{E}\varepsilon_{4}=\mathbb{E}\sum_{i,j}\frac{\jmath u^{3}}{3M}\frac{\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}]_{j,i}={\mathcal{O}}(\frac{u^{3}\mathcal{P}(\frac{1}{z})(1+\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{5}{2}}})}{M^{2}z})={\mathcal{O}}(\frac{u^{3}\mathcal{P}(\frac{1}{z})}{zM^{2}}).$ (109)
@@
$\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ $=M\omega^{2}_{z}\overline{\omega}^{2}_{z}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\overline{\omega}^{2}_{z}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{\jmath u\overline{\omega}_{z}\sqrt{n}}{\sqrt{M^{3}}}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ (110)
$+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{z}f({\mathbb{A}},u)).$
@@
$f({\mathbb{A}},u)=(\frac{1}{M}+\frac{u\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}+\frac{u^{2}(M+\sqrt{\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}})}{M^{2}}+\frac{u^{3}(M+\sqrt{\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}})}{M^{3}}).$ (111)
@@
Now we turn to evaluate $\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ in the second last line of (C-B1). By the integration by parts formula, we have
$\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=\sum_{i,j}Z_{j,i}^{*}[{\mathbb{Q}}(z){\mathbb{Z}}]_{j,i}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ (112)
$=\mathbb{E}[\operatorname{Tr}{\mathbb{Q}}(z)-\frac{1}{L}\operatorname{Tr}{\mathbb{Q}}(z)\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}+\frac{\jmath u\sqrt{n}}{\sqrt{M}L}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}-\frac{\jmath un}{\sqrt{Mn}L}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}$
@@
$+\frac{\jmath un}{\sqrt{Mn}L}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}-\frac{u^{2}\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Z_{j,i}}\frac{[{\mathbb{Q}}(z){\mathbb{Z}}]_{j,i}}{2L}+\frac{\jmath u^{3}\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Z_{j,i}}\frac{[{\mathbb{Q}}(z){\mathbb{Z}}]_{j,i}}{3L}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}.$
@@
By the analysis of (103), we can show that $\frac{1}{L}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}$ and $\frac{1}{L}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}$ are ${\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{z^{2}M^{\frac{5}{2}}})$. By following the same procedure from (103) to (106), we can show that the last two terms in (112) are ${\mathcal{O}}(\frac{1}{Mz})$ and ${\mathcal{O}}(\frac{u^{3}\mathcal{P}(\frac{1}{z})}{zM^{2}})$, respectively. Therefore, we have
$\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z)\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}(1-\frac{M}{L}\omega_{z}\overline{\omega}_{z})-\delta_{z}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+M\delta_{z}\omega_{z}\overline{\omega}_{z}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ (113)
$+\frac{\jmath u\sqrt{n}}{\sqrt{M}L}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(f({\mathbb{A}},u,z)).$

### 09 Apr 2024 v5

@@ C-B1 The evaluation of $U_{1}$
$\mathbb{E}[\varepsilon_{1}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=\mathbb{E}[\varepsilon_{1}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\mathrm{Cov}(\varepsilon_{1},\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}),$ (102)
and $\mathrm{Cov}(\varepsilon_{1},\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}})\leq\mathrm{Var}^{\frac{1}{2}}(\varepsilon_{1})\mathrm{Var}^{\frac{1}{2}}(\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}})={\mathcal{O}}(\frac{u\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}\mathcal{P}_{2}(\frac{1}{z})}{zM^{\frac{3}{2}}})$. By the integration by parts formula, we have
$\mathbb{E}\varepsilon_{1}=-\frac{\jmath un}{\sqrt{M^{3}n}}\operatorname{Tr}({\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}})=-\frac{\jmath un}{\sqrt{M^{3}n}}\Bigl{(}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{A}}\mathbb{E}\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}})]}{M}$ (103)
$-\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})]}{M}-\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})\operatorname{Tr}({\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}})]}{M}\Bigr{)}$
@@
Noticing that $\operatorname{Tr}({\mathbb{A}})=\operatorname{Tr}({\mathbb{I}}_{M}-\frac{{\mathbb{X}}{\mathbb{X}}^{H}}{n})=0$, we have $E_{1}=0$. By (60) and $\operatorname{Tr}({\mathbb{A}}^{2})={\mathcal{O}}(M^{2})$, we have
$|E_{2}|=\frac{un}{\sqrt{M^{3}n}}\Bigl{|}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})]}{M}+\mathrm{Cov}\Bigl{(}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)),\frac{\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})}{M}\Bigr{)}\Bigr{|}$ (104)
$\leq\frac{Kun}{\sqrt{M^{3}n}}\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{z^{2}M^{\frac{3}{2}}}={\mathcal{O}}\Bigl{(}\frac{u\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{z^{2}n^{\frac{5}{2}}}\Bigr{)}.$
@@
$|E_{3}|\leq\frac{un}{\sqrt{M^{3}n}}\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{z^{2}M}={\mathcal{O}}\Bigl{(}\frac{u\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{z^{2}n^{\frac{5}{2}}}\Bigr{)}.$ (105)
A similar analysis can be performed to obtain $\mathbb{E}[\varepsilon_{2}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]={\mathcal{O}}(\frac{u\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{z^{2}n^{\frac{5}{2}}})$ and $\mathbb{E}[\varepsilon_{3}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=\mathbb{E}[\varepsilon_{3}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}_{3}(\frac{1}{z})(1+\sqrt{\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}})}{M^{2}z^{2}}\Bigr{)}$ in the last line of (C-B1). Now we turn to evaluate $\mathbb{E}\varepsilon_{3}$, which is performed as
$\mathbb{E}[\varepsilon_{3}]=-\mathbb{E}[\sum_{i,j}\frac{u^{2}}{2M}\frac{\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}]_{j,i}]=-\frac{u^{2}n}{Mn}[\frac{1}{M}\mathbb{E}[\operatorname{Tr}({\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]$ (106)
$-\frac{1}{M}\mathbb{E}[\operatorname{Tr}({\mathbb{H}}{\mathbb{H}}^{H}({\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})^{2}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]+\frac{\sigma^{2}}{2M}\mathbb{E}[\operatorname{Tr}({\mathbb{X}}{\mathbb{X}}^{H}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}})]$
$-\frac{\sigma^{2}\mathbb{E}[\operatorname{Tr}({\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{X}}{\mathbb{X}}^{H}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}-\frac{\sigma^{2}\mathbb{E}[\operatorname{Tr}({\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{X}}{\mathbb{X}}^{H}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}$
@@
Similarly, we can show that $E_{3,i},i=2,...,3$ are all ${\mathcal{O}}(\frac{1}{Mz})$ terms. We can evaluate $E_{3,4}$ as
$\frac{2nu^{2}}{Mn}E_{3,4}$ $=-\frac{2nu^{2}}{Mn}\frac{\sigma^{2}\mathbb{E}[\operatorname{Tr}({\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{X}}{\mathbb{X}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{Mn}=\frac{2nu^{2}}{Mn}[\frac{\sigma^{2}\mathbb{E}[\operatorname{Tr}({\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}$ (108)
$-\frac{\sigma^{2}\mathbb{E}[\operatorname{Tr}({\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}]={\mathcal{O}}\Biggl{(}\frac{\mathcal{P}(\frac{1}{z})u^{2}(1+\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{5}{2}}})}{zM}\Biggr{)}.$
Therefore, $\mathbb{E}[\varepsilon_{3}]={\mathcal{O}}\Bigl{(}\frac{u^{2}\mathcal{P}(\frac{1}{z})(1+\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{5}{2}}})}{Mz}\Bigr{)}={\mathcal{O}}(\frac{u^{2}\mathcal{P}(\frac{1}{z})}{zM})$. According to the derivation in (106) to (108), we can also show that
$\mathbb{E}[\varepsilon_{4}]=\mathbb{E}[\sum_{i,j}\frac{\jmath u^{3}}{3M}\frac{\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}]_{j,i}]={\mathcal{O}}\Biggl{(}\frac{u^{3}\mathcal{P}(\frac{1}{z})(1+\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{5}{2}}})}{M^{2}z}\Biggr{)}={\mathcal{O}}\Bigl{(}\frac{u^{3}\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)}.$ (109)
@@
$\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}})]$ $=M\omega^{2}_{z}\overline{\omega}^{2}_{z}\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\overline{\omega}^{2}_{z}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\frac{\jmath u\overline{\omega}_{z}\sqrt{n}}{\sqrt{M^{3}}}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ (110)
$+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{z}f({\mathbb{A}},u)\Bigr{)}.$
@@
$f({\mathbb{A}},u)=\Biggl{(}\frac{1}{M}+\frac{u\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{3}{2}}}+\frac{u^{2}(M+\sqrt{\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}})}{M^{2}}+\frac{u^{3}(M+\sqrt{\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}})}{M^{3}}\Biggr{)}.$ (111)
@@
Now we turn to evaluate $\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ in the second last line of (C-B1). By the integration by parts formula, we have
$\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=\sum_{i,j}\mathbb{E}[Z_{j,i}^{*}[{\mathbb{Q}}(z){\mathbb{Z}}]_{j,i}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ (112)
$=\mathbb{E}\Bigl{[}[\operatorname{Tr}({\mathbb{Q}}(z))-\frac{1}{L}\operatorname{Tr}({\mathbb{Q}}(z)\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})+\frac{\jmath u\sqrt{n}}{\sqrt{M}L}\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}})-\frac{\jmath un}{\sqrt{Mn}L}\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}})$
@@
$+\frac{\jmath un}{\sqrt{Mn}L}\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}})-\frac{u^{2}\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Z_{j,i}}\frac{[{\mathbb{Q}}(z){\mathbb{Z}}]_{j,i}}{2L}+\frac{\jmath u^{3}\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Z_{j,i}}\frac{[{\mathbb{Q}}(z){\mathbb{Z}}]_{j,i}}{3L}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}\Bigr{]}.$
@@
By the analysis of (103), we can show that $\frac{1}{L}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}})]$ and $\frac{1}{L}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}})]$ are ${\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{z^{2}M^{\frac{5}{2}}})$. By following the same procedure from (103) to (106), we can show that the last two terms in (112) are ${\mathcal{O}}(\frac{1}{Mz})$ and ${\mathcal{O}}(\frac{u^{3}\mathcal{P}(\frac{1}{z})}{zM^{2}})$, respectively. Therefore, we have
$\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z))\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}](1-\frac{M}{L}\omega_{z}\overline{\omega}_{z})-\delta_{z}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+M\delta_{z}\omega_{z}\overline{\omega}_{z}\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ (113)
$+\frac{\jmath u\sqrt{n}}{\sqrt{M}L}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}(f({\mathbb{A}},u,z)),$